Snowshoe Mountain Excursions (SMEx) had a share price at the start of the year of $96.12 and paid a dividend of $2.14 at the end of the year. If SMEx has an equity cost of capital of 14%, what price would you expect to pay for a share of stock just after it pays its dividend at the end of the year?

A) $86.19

B) $109.58

C) $107.44

Please show all work.
Answer #1

As the cost of capital is 14%, so if it did not give any dividend,the price of share would be (cum dividend)= 96.12*1.14=$109.58 at the end of year

As it has given dividend of $2.14, the price of share=109.58-2.14=$107.44 (ex-dividend price)

Option C is correct.